A utility pole is 8 m high. A cable is stretched from the top of the pole to a point in the ground that is 6 m from the bottom of the pole. How long is the cable?

Answer by math_tutor2020(3817)


Assuming the pole is perfectly vertical, and the ground is completely horizontal, then a right triangle forms.

a = 8 = vertical leg
b = 6 = horizontal leg
c = unknown hypotenuse

Apply the pythagorean theorem
a^2 + b^2 = c^2
c = sqrt(a^2 + b^2)
c = sqrt(8^2 + 6^2)
c = sqrt(64+36)
c = sqrt(100)
c = 10

Answer: The cable is 10 meters long.